The sum of magnitudes of two forces acting at a point 16 N.the resultant of these forces is perpendicular to the smaller force of 8N.if the smaller force magnitude x, the value of x is 

Grade:11

1 Answers

Vikas TU
6 years ago

Now we know that abs(a)+abs(b)=16abs(a)+abs(b)=16

And since this is a vector triangle, using Pythagoras theorem. a2b2=64a2−b2=64

From abs(a)+abs(b)=16abs(a)+abs(b)=16

Make abs(b)abs(b) the subject followed by squaring the equation

This will give you

b2=25632abs(a)+a2b2=256−32abs(a)+a2

(abs(x) squared= x^2 due to commutative law of vector scalar multiplication)

Substitute into the Pythagoras theorem equation and you will get

32abs(a)=32032abs(a)=320

abs(a)=10abs(a)=10

And from there substitute back into equation to get abs(b)=6